OpenStudy (anonymous):

Factoring csc^2x - 1 ? I have it to (csc x + 1)(csc x -1) but I don't know what to do now.

OpenStudy (anonymous):

what are you trying to do?

OpenStudy (anonymous):

just factor?

OpenStudy (anonymous):

It says I must convert it to one single trig function.

OpenStudy (amistre64):

consider the most fundamental trig identity: sin^2 + cos^2 = 1

OpenStudy (amistre64):

since csc = 1/sin .... divide it all by sin^2

OpenStudy (anonymous):

But how can I make the +1 / -1 disappear?

OpenStudy (amistre64):

sin^2 + cos^2 = 1 sin^2/sin^2 + cos^2/sin^2 = 1/sin^2 1 + cot^2 = csc^2 cot^2 = csc^2 - 1

OpenStudy (amistre64):

if you go the factoring route .... csc = sqrt(1+cot) i spose

OpenStudy (anonymous):

Thank you very much. I understand it very well now.

OpenStudy (amistre64):

good